# Stabilization of intersection Betti numbers for moduli spaces of one-dimensional sheaves on surfaces

### v1

@@ 3.2. Local versality and smoothness in high codimension
Let $\ell\geq 2$ be an integer. If $\beta$ is $(2\ell-2)$-very ample, then there exists an open subscheme $V_{\ell}\subset|\beta|$ such that
@@
    Every point of $V_{\ell}$ corresponds to a reduced curve;
@@
    $\operatorname{codim}(|\beta|\setminus V_{\ell},|\beta|)\geq\ell+1$;
@@
    The universal family $\pi|_{\pi^{-1}(V_{\ell})}:\pi^{-1}(V_{\ell})\to V_{\ell}$ of curves in $V_{\ell}$ is locally versal.
@@
We use an argument in [16] to control the locus of curves that are not $\ell$-nodal. Here, a curve is called $\ell$-nodal if it is reduced and has no singularities other than $\ell$ nodes. Since $\beta$ is also $\ell$-very ample, by the proof of [16, Proposition 2.1], there is an open subscheme $V_{\ell}\subset|\beta|$ such that
  • $\operatorname{codim}(|\beta|\setminus V_{\ell},|\beta|)>\ell$.
  • Curves in $V_{\ell}$ are either $\ell$-nodal, or reduced of geometric genus $p_{a}(\beta)-\ell$.
Note that for a curve $C\in V_{\ell}$ which is not $\ell$-nodal, its geometric genus is
(3.8) $p_{a}(\beta)-\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)>p_{a}(\beta)-\ell.$
By (3.8) and Milnor’s formula (3.6), we have
$\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)\leq\sum_{p\in C_{\mathrm{sing}}}\mu_{p}(C)\leq 2\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)<2\ell.$
It remains to show Property (3). For any curve $D\in V_{\ell}$, the map $k_{\pi,D}^{\text{loc}}$ in (3.7) can be identified with
$\bar{r}_{W}:H^{0}({\mathcal{O}}_{S}(\beta))/\langle s_{D}\rangle\to H^{0}({\mathcal{O}}_{S}(\beta)|_{W})$
induced by the restriction to a 0-dimensional closed subscheme $W\subset S$ of length
$\sum_{p\in D_{\mathrm{sing}}}\tau_{p}(D)<2\ell$
by Lemma 3.3, where $s_{D}\in H^{0}({\mathcal{O}}_{S}(\beta))$ is a defining section of $D$. Hence, $\bar{r}_{W}$ is surjective by the ($2\ell-2$)-very ampleness of $\beta$. ∎
Combining the above proposition with Theorem 3.5, we can prove Theorem 1.2.
@@
Let $V_{\ell}$ be as in Proposition 3.6. Take $U_{\ell}=V_{\ell}\cap|\beta|^{\mathrm{int}}$. By Property (2) of $V_{\ell}$ in Proposition 3.6, we have
$\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{\ell},|\beta|^{\mathrm{int}})\geq\ell+1.$
It remains to show that $h^{-1}(U_{\ell})$ is smooth. Note that $h^{-1}(U_{\ell})$ is the relative compactified Jacobian of degree $p_{a}(\beta)-1+\chi$ associated with the universal family of curves in $U_{\ell}$. This universal family is locally versal by Property (3) of $V_{\ell}$ in Proposition 3.6. Thus, the smoothness of $h^{-1}(U_{\ell})$ follows from Theorem 3.5. ∎
@@ 4. Intersection Betti numbers
In this section, we generalize the approach in [29] to prove Theorem 1.1. For any open subscheme $U\subset|\beta|$, let
@@
First, we focus on a smooth open subscheme $h^{-1}(U)$ of $M_{\beta,\chi}$ and prove the following result concerning the dimensions of graded pieces of the perverse filtration induced by the restriction of $h:M_{\beta,\chi}\to|\beta|$ (see (2.4)).
@@
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $h_{U}:h^{-1}(U)\to U$ be the restriction of $h$. For given $i,j\in\mathbb{Z}_{\geq 0}$, if $i+j\leq N(U)$, the divisor $\beta$ is (i1)(i-1)-very ample, and $2\dim|\beta|\geq 3i+j$, then
(4.9) $n_{h_{U}}^{i,j}=n_{\infty}^{i,j},$
where $n_{\infty}^{i,j}$ is the coefficient of $q^{i}t^{j}$ in the expansion of
@@
Recall Göttsche’s formula [8] for $\sum_{n\geq 0}\sum_{i\geq 0}b_{i}(S^{[n]})z^{i}w^{n}$:
@@
the product formula $H(q,t)$ in (4.10) satisfies the following equality
@@
Let $k\in\mathbb{Z}_{\geq 0}$. In the setting of Theorem 4.1, if $k\leq\min\{N(U),2\dim|\beta|/3\}$ and $\beta$ is (k1k-1)-very ample, then
@@
4.1. Strategy of the proof
The proof of Theorem 4.1 consists of the following steps.
Step 1. Relate $h^{-1}(U)$ to the relative Hilbert scheme ${\mathcal{C}}^{[n]}_{U}$. By (2.5), for any $\ell,m\in\mathbb{Z}_{\geq 0}$,
(4.13) $b_{m}({\mathcal{C}}^{[\ell]}_{U})=\sum_{i+j\leq\ell;\,i,j\geq 0}n_{h_{U}}^{i,m-i-2j}.$
Step 2. Calculate the Betti numbers of ${\mathcal{C}}^{[\ell]}_{U}$ using Proposition 2.3.
@@ 4.1. Strategy of the proof
In the setting of Theorem 4.1, if $m\leq\min\{N(U),2\dim|\beta|-2\ell\}$ and $\beta$ is ($\ell-1$)-very ample, then
(4.14) $b_{m}({\mathcal{C}}^{[\ell]}_{U})=\sum_{n=0}^{\lfloor m/2\rfloor}b_{m-2n}(S^{[\ell]}).$
@@
Since $\pi^{[\ell]}:{\mathcal{C}}^{[\ell]}\to|\beta|$ has fibers of dimension $\ell$ by Proposition 2.1, we have
$\operatorname{codim}({\mathcal{C}}^{[\ell]}\setminus{\mathcal{C}}_{U}^{[\ell]},{\mathcal{C}}^{[\ell]})=\operatorname{codim}(|\beta|\setminus U,|\beta|).$
By Proposition 2.3, the $(\ell-1)$-very ampleness of $\beta$ implies that $\mathcal{C}$$^{[\ell]}$ is a projective bundle over $S^{[\ell]}$. In particular, $\mathcal{C}$$^{[\ell]}$ is irreducible and smooth. Then it follows from [29, Lemma 2.4] (or the proof of Lemma 5.6) that for $m\leq N(U)$,
$b_{m}({\mathcal{C}}^{[\ell]}_{U})=b_{m}({\mathcal{C}}^{[\ell]}).$
Combining with
$\dim{\mathcal{C}}^{[\ell]}-\dim S^{[\ell]}=\dim|\beta|-\ell\geq m/2$
by our assumption, we conclude
$b_{m}({\mathcal{C}}^{[\ell]}_{U})=b_{m}({\mathcal{C}}^{[\ell]})=b_{m}(S^{[\ell]}\times\mathbb{P}^{\dim|\beta|-\ell})=\sum_{n=0}^{\lfloor m/2\rfloor}b_{m-2n}(S^{[\ell]}),$
which proves (4.14). ∎
Since all Betti numbers of $S^{[\ell]}$ are determined by Göttsche’s formula (4.11), we can prove Theorem 4.1 by induction on $i$, combining (4.13) and (4.14).
@@
We proceed by induction on $i$ to prove (4.9). For $i=0$ and $j\leq\min\{N(U),2\dim|\beta|\}$,
$n_{h_{U}}^{0,j}$ $=\dim P_{0}H^{j}(h^{-1}(U))$
@@
which proves the induction base of (4.9). Suppose the result is true for $(i^{\prime},j^{\prime})$ with $i^{\prime}<i$ and $i^{\prime}+j^{\prime}\leq i+j$. For $m\leq i+j$, it follows by (4.13), our induction hypothesis, and (4.14) that
$n_{h_{U}}^{i,m-i}$ $=\sum_{i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},m-i^{\prime}-2j^{\prime}}-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
$=\left(b_{m}({\mathcal{C}}^{[i]}_{U})-b_{m}({\mathcal{C}}^{[i-1]}_{U})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{\infty}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
$=\sum_{n=0}^{\lfloor m/2\rfloor}\left(b_{m-2n}(S^{[i]})-b_{m-2n}(S^{[i-1]})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{\infty}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
$=n_{\infty}^{i,m-i},$
where the last identity is essentially the same as [29, (31)] by Göttsche’s formula [8] (see also Remark 4.2). ∎
Finally, we prove our first main theorem stated in the introduction.
@@
By the upper semicontinuity of fiber dimensions, we have
@@
Since $\beta$ is $\max\{2,2k-2\}$-very ample, let $U_{k}\subset|\beta|^{\mathrm{int}}$ be as in Theorem 1.2 (set $U_{k}\vcentcolon=U_{2}$ if $k\leq 1$). Using (4.15), the assumption $k\leq c-1$, and Theorem 1.2, we obtain the following properties:
@@
By Corollary 4.3, we have
@@
Then by [23, Theorem 6.7.4] and the smoothness of $h^{-1}(U_{k})$, we have isomorphisms
$\operatorname{IH}^{k}(M_{\beta,\chi})\cong\operatorname{IH}^{k}(h^{-1}(U_{k}))\cong H^{k}(h^{-1}(U_{k})),$
and therefore
$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{\beta,\chi})=b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty}$
by (4.16), which completes the proof. ∎

### v2

@@ 3.2. Local versality and smoothness in high codimension
If $\beta$ is $\max\{1,2k-2\}$-very ample, then there exists an open subscheme $V_{k}\subset|\beta|$ such that
@@
    every point of $V_{k}$ corresponds to a reduced curve;
@@
    $\operatorname{codim}(|\beta|\setminus V_{k},|\beta|)\geq k+1$;
@@
    the universal family $\pi|_{\pi^{-1}(V_{k})}:\pi^{-1}(V_{k})\to V_{k}$ of curves in $V_{k}$ is locally versal.
@@
By setting $V_{0}\vcentcolon=V_{1}$, we may assume $k>0$. The assumption on $\beta$ implies that $\beta$ is both $k$-very ample and (2k22k-2)-very ample. We use an argument in [22] to control the locus of curves that are not $k$-nodal. Here, a curve is called $k$-nodal if it is reduced and has no singularities other than $k$ nodes. Since $\beta$ is $k$-very ample, by the proof of [22, Proposition 2.1], there is an open subscheme $V_{k}\subset|\beta|$ satisfying the following conditions:
  • $\operatorname{codim}(|\beta|\setminus V_{k},|\beta|)>k$.
  • Curves in $V_{k}$ are either $k$-nodal, or reduced of geometric genus $p_{a}(\beta)-k$.
Note that for a curve $C\in V_{k}$ which is not $k$-nodal, its geometric genus $p_{g}(C)$ satisfies
$p_{g}(C)=p_{a}(\beta)-\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)>p_{a}(\beta)-k.$
From this and Milnor’s formula (3.7), we obtain
(3.9) $\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)\leq\sum_{p\in C_{\mathrm{sing}}}\mu_{p}(C)\leq 2\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)<2k.$
It remains to show Property (3). For any curve $D\in V_{k}$, let $s_{D}\in H^{0}(S,{\mathcal{O}}_{S}(\beta))$ be a defining section of $D$. The map $\kappa_{\pi,D}^{\text{loc}}$ as in (3.8) can be identified with the map
$\bar{r}_{Z_{D}}:H^{0}(S,{\mathcal{O}}_{S}(\beta))/\langle s_{D}\rangle\to H^{0}(Z_{D},{\mathcal{O}}_{S}(\beta)|_{Z_{D}})$
induced by restriction to a 0-dimensional closed subscheme $Z_{D}\subset S$. By (3.9) (which still holds for a $k$-nodal curve) and Lemma 3.3, the length $h^{0}({\mathcal{O}}_{Z_{D}})$ of $Z_{D}$ satisfies
$h^{0}({\mathcal{O}}_{Z_{D}})=h^{0}({\mathcal{E}xt}^{1}(\Omega_{D},{\mathcal{O}}_{D}))=\sum_{p\in D_{\mathrm{sing}}}\tau_{p}(D)<2k.$
Hence, $\kappa_{\pi,D}^{\text{loc}}$ is surjective by the (2k22k-2)-very ampleness of $\beta$. ∎
Combining the above proposition with Proposition 3.5, we can prove Theorem 1.2.
@@
Let $V_{k}$ be as in Proposition 3.6. Take $U_{k}=V_{k}\cap|\beta|^{\mathrm{int}}$. By Property (2) of $V_{k}$ in Proposition 3.6, we have
$\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1.$
It remains to show that $h^{-1}(U_{k})$ is smooth. Note that $h^{-1}(U_{k})$ is the relative compactified Jacobian of degree $p_{a}(\beta)-1+\chi$ associated with the universal family of curves in $U_{k}$. This universal family is locally versal by Property (3) of $V_{k}$ in Proposition 3.6. Thus, the smoothness of $h^{-1}(U_{k})$ follows from Proposition 3.5. ∎
@@ 4. Intersection Betti numbers
In this section, we generalize the approach in [36] to prove Theorem 1.1. To start with, we review some properties of relative Hilbert schemes of points over $|\beta|$, which will play an important role in the proof.
4.1. Relative Hilbert schemes of points
Let $\pi:{\mathcal{C}}\to|\beta|$ be the universal family of curves in $|\beta|$. For $k\in\mathbb{Z}_{\geq 0}$, denote by
$\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$
the relative Hilbert scheme of $k$ points on the fibers of $\pi$. For $C\in|\beta|$, the fiber of $\pi^{[k]}$ over $C$ is the Hilbert scheme $C^{[k]}$ parametrizing 0-dimensional, length $k$ closed subschemes of $C$. Note that $\pi^{[0]}:|\beta|\to|\beta|$ is the identity and $\pi=\pi^{[1]}:{\mathcal{C}}\to|\beta|$. The following property of $\pi^{[k]}$ is a corollary of [24, Theorem 1.1] (see also [42, Proposition 2.7]).
Lemma 4.1.
The morphism $\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$ has fibers of the same dimension $k$.
There is a natural morphism $\sigma_{k}:{\mathcal{C}}^{[k+1]}\to S^{[k+1]}$ defined by
$Z\subset C]\mapsto[Z\subset S],$
where $C\in|\beta|$, and $Z\subset C$ is a 0-dimensional closed subscheme of length $k+1$. The fiber of $\sigma_{k}$ over $Z\subset S]\in S^{[k+1]}$ is the projective space $\mathbb{P}(H^{0}(S,{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))^{\vee})$, where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$. A well-known connection between $k$-very ampleness and the structure of ${\mathcal{C}}^{[k+1]}$ as a scheme over $S^{[k+1]}$ via $\sigma_{k}$ is as follows (e.g., [42, Proposition 2.6]).
Lemma 4.2.
Recall that for the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$, the fibers of $h$ over the locus $|\beta|^{\mathrm{int}}$ of integral curves are the compactified Jacobians of degree $p_{a}(\beta)-1+\chi$ and have dimension $p_{a}(\beta)$. For any open subscheme $U\subset|\beta|$, let $h_{U}:h^{-1}(U)\to U$ be the restriction of $h$, and write ${\mathcal{C}}_{U}^{[k]}$ for $(\pi^{[k]})^{-1}(U)$. Later, we will use the following result, due to Maulik–Yun [28, Theorem 2.13] and Migliorini–Shende [31, Corollary 2].
Theorem 4.3 ([28, 31] [^2] The results of [28, 31] apply since the smoothness of $h^{-1}(U)$ implies, via a section of the universal curve (étale locally on $U$, avoiding the singularities of fibers), that the relative compactified Jacobians $\overline{J}_{U}^{e}$ for all $e\in\mathbb{Z}$ are smooth, and by [41, Proposition 14], so are the relative Hilbert schemes ${\mathcal{C}}^{[k]}_{U}$ for all $k\in\mathbb{Z}_{\geq 0}$.).
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $P_{\bullet}H^{*}(h^{-1}(U))$ be the perverse filtration induced by $h_{U}:h^{-1}(U)\to U$. For any $k,\ell\in\mathbb{Z}_{\geq 0}$, we have an isomorphism of vector spaces
(4.10) $H^{\ell}\left({\mathcal{C}}^{[k]}_{U}\right)\cong\bigoplus_{i+j\leq k;\,i,j\geq 0}\mathrm{Gr}_{i}^{P}H^{\ell-2j}\left(h^{-1}(U)\right).$
Remark 4.4.
When $k=0$, the isomorphism (4.10) reads
$H^{\ell}(U)\cong\mathrm{Gr}_{0}^{P}H^{\ell}\left(h^{-1}(U)\right),$
which can be obtained as a consequence of Ngô’s support theorem [33, Théorème 7.2.1] (see [28, Theorem 2.4, §2.5] or [31, Corollary 9]). Ngô’s support theorem implies that $h_{U}:h^{-1}(U)\to U$ has full support. More precisely, there is an isomorphism
$Rh_{U*}{\mathbb{Q}}_{h^{-1}(U)}[\dim U]\simeq\bigoplus_{i=0}^{2p_{a}(\beta)}\operatorname{IC}(\wedge^{i}R^{1}\pi^{s}_{*}{\mathbb{Q}}_{{\mathcal{C}}_{U}^{\mathrm{sm}}})[-i],$
where $\pi^{s}:{\mathcal{C}}_{U}^{\mathrm{sm}}\to U^{\mathrm{sm}}$ is the universal family of smooth curves in $U$.
4.2. Strategy of the proof
For any open subscheme $U\subset|\beta|$, let
@@ 4.2. Strategy of the proof
As a preliminary step, we focus on a smooth open subscheme $h^{-1}(U)$ of $M_{\beta,\chi}$ and prove the following result concerning graded dimensions of the perverse filtration induced by the restriction $h_{U}:h^{-1}(U)\to U$ of $h:M_{\beta,\chi}\to|\beta|$ (see (2.6)).
@@
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth. For any $i,j\in\mathbb{Z}_{\geq 0}$ satisfying $i+j\leq N(U)$ and $3i+j\leq 2\dim|\beta|$, if $\beta$ is (i1)(i-1)-very ample, then
$n_{h_{U}}^{i,j}=[H(q,t)]^{i,j},$
where $H(q,t)]^{i,j}$ is the coefficient of $q^{i}t^{j}$ in the expansion of
@@
Recall Göttsche’s formula [11] for $\sum_{k\geq 0}\sum_{i\geq 0}b_{i}(S^{[k]})z^{i}w^{k}$:
@@
the product formula $H(q,t)$ in (4.11) satisfies the following equality:
@@
In the setting of Theorem 4.5, if $k\leq\min\{N(U),2\dim|\beta|/3\}$ and $\beta$ is (k1k-1)-very ample, then
@@
We prove Theorem 4.5 by relating $h^{-1}(U)$ to the relative Hilbert schemes ${\mathcal{C}}^{[k]}_{U}$. By (4.10), for any $k,\ell\in\mathbb{Z}_{\geq 0}$,
(4.14) $b_{\ell}({\mathcal{C}}^{[k]}_{U})=\sum_{i+j\leq k;\,i,j\geq 0}n_{h_{U}}^{i,\ell-i-2j}.$
To calculate the Betti numbers of ${\mathcal{C}}^{[k]}_{U}$, we use Lemmas 4.1 and 4.2.
@@
In the setting of Theorem 4.5, if $\ell\leq\min\{N(U),2\dim|\beta|-2k\}$ and $\beta$ is (k1k-1)-very ample, then
(4.15) $b_{\ell}({\mathcal{C}}^{[k]}_{U})=\sum_{r=0}^{\lfloor\ell/2\rfloor}b_{\ell-2r}(S^{[k]}).$
@@
Since all fibers of $\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$ have dimension $k$ by Lemma 4.1, we have
$\operatorname{codim}({\mathcal{C}}^{[k]}\setminus{\mathcal{C}}_{U}^{[k]},{\mathcal{C}}^{[k]})=\operatorname{codim}(|\beta|\setminus U,|\beta|).$
By Lemma 4.2, the (k1)(k-1)-very ampleness of $\beta$ implies that ${\mathcal{C}}^{[k]}$ is a projective bundle over $S^{[k]}$. In particular, ${\mathcal{C}}^{[k]}$ is irreducible and smooth. Then it follows from [36, Lemma 2.4] (or the proof of Lemma 5.3) that for $\ell\leq N(U)$,
$b_{\ell}({\mathcal{C}}^{[k]}_{U})=b_{\ell}({\mathcal{C}}^{[k]}).$
Combining this with
$\dim{\mathcal{C}}^{[k]}-\dim S^{[k]}=\dim|\beta|-k\geq\ell/2,$
which follows from our assumption, we conclude that
$b_{\ell}({\mathcal{C}}^{[k]}_{U})=b_{\ell}({\mathcal{C}}^{[k]})=b_{\ell}(S^{[k]}\times\mathbb{P}^{\dim|\beta|-k})=\sum_{r=0}^{\lfloor\ell/2\rfloor}b_{\ell-2r}(S^{[k]}),$
thereby completing the proof of (4.15). ∎
Since all Betti numbers of $S^{[k]}$ are determined by Göttsche’s formula (4.12), now we can prove Theorem 4.5 by combining (4.14) and (4.15).
@@
We use induction on $i$. For $i=0$ and $j\leq\min\{N(U),2\dim|\beta|\}$,
$n_{h_{U}}^{0,j}$ $=\dim_{\mathbb{Q}}\mathrm{Gr}_{0}^{P}H^{j}(h^{-1}(U))$
@@
which proves the induction base. Then suppose that $n_{h_{U}}^{i^{\prime},j^{\prime}}=[H(q,t)]^{i^{\prime},j^{\prime}}$ holds for any $i^{\prime},j^{\prime}\in\mathbb{Z}_{\geq 0}$ satisfying $i^{\prime}<i$ and $i^{\prime}+j^{\prime}\leq i+j$. For any $\ell\leq i+j$, it follows from (4.14), our induction hypothesis, and (4.15) that
$n_{h_{U}}^{i,\ell-i}$ $=\sum_{i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
$=\left(b_{\ell}({\mathcal{C}}^{[i]}_{U})-b_{\ell}({\mathcal{C}}^{[i-1]}_{U})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}[H(q,t)]^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
$=\sum_{r=0}^{\lfloor\ell/2\rfloor}\left(b_{\ell-2r}(S^{[i]})-b_{\ell-2r}(S^{[i-1]})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}[H(q,t)]^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
$=[H(q,t)]^{i,\ell-i},$
where the last identity is essentially equivalent to [36, (31)] and follows from Remark 4.6. This completes the induction step and hence the proof. ∎
Finally, we prove our first main theorem stated in Section 1.
@@
By the upper semicontinuity of fiber dimension, we have
@@
Since $\beta$ is $\max\{1,2k-2\}$-very ample, let $U_{k}\subset|\beta|^{\mathrm{int}}$ be as in Theorem 1.2. Using (4.16), the assumption $k\leq c-1$, and Theorem 1.2, we obtain the following properties:
  1. (1)
@@
Note that $|\beta|$ contains a connected smooth curve by Bertini’s theorem. Applying Corollary 4.7 with $U=U_{k}$, we have
@@
Then by [30, Theorem 6.7.4] and the smoothness of $h^{-1}(U_{k})$, we have isomorphisms
$I\!H^{k}(M_{\beta,\chi})\cong I\!H^{k}(h^{-1}(U_{k}))\cong H^{k}(h^{-1}(U_{k})),$
and therefore, by (4.17),
$\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})=b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty},$
which completes the proof. ∎